An object with an initial velocity of 4 m/s moves along a straight axis under constant acceleration. 3 seconds later, its velocity is 14m/s. How far did it travel during this period
1
Expert's answer
2019-11-21T10:02:30-0500

The distance traveled by an object


"d=\frac{v_f+v_i}{2}t=\frac{(14\:\rm m/s)+(4\:\rm m/s)}{2}\times 3\:\rm s=27\: m"